# Efficient Nearest Neighbor Search Using Dynamic Programming

### v3

@@ Abstract
When dealing with point clouds distributed on manifold surfaces in 3D space, or when the query point is far from the data, the efficiency of traditional nearest neighbor search algorithms (e.g., KD Tree and R Tree) may degrade. In extreme cases, the complexity of the query can approach $O(n)$. In this paper, we propose a novel dynamic programming technique that precomputes a Directed Acyclic Graph (DAG) to enable more efficient nearest neighbor queries for 2D manifold data. By leveraging this structure, only a small number of distance comparisons between point pairs are required to accurately identify the nearest neighbor. Extensive experimental results demonstrate that our method achieves query speeds that are 1x-10x faster than traditional methods. Moreover, our algorithm exhibits significant potential. It achieves query efficiency comparable to KD-trees on uniformly distributed point clouds. Additionally, our algorithm supports nearest neighbor queries among the first $k$ points. Coupled with our algorithm, a farthest point sampling algorithm with lower complexity can also be implemented. Furthermore, our method has the potential to support nearest neighbor queries with different types of primitives and distance metrics. We believe that the method proposed in this paper represents the most concise and straightforward exact nearest neighbor search algorithm currently available, and it will contribute significantly to advancements in the field.
@@ I Introduction
As shown in the inset figure, we have two point clouds with distinct distributions: one uniformly distributed within a unit cube, and the other distributed across the surface of the Earth. Additionally, we generate two sets of query points—one within the bounding box of the input data, and the other within 8x the bounding box. It can be observed that the query performance of both KD Tree and R Tree significantly degrades under these conditions.
@@
  • If the point cloud possesses an order, our method can facilitate the identification of the nearest point to $q$ among $\{p_{i}\}_{i=1}^{m}$, for any $m$, where $m<n$. This approach can be employed to reduce the time complexity of Density Peak Clustering [9], thus enhancing its applicability in large-scale scenarios.
@@
Additionally, our algorithm exhibits significant potential. For instance, under various primitive types and distance metrics, as long as the incremental Voronoi or Delaunay construction algorithms for the primitives, as well as the distance query function from a point in space to the primitive, are provided, our nearest neighbor search framework can support nearest neighbor searches in these contexts. However, further development will be necessary in future work.
II Releated Work
@@ II-A Exact Nearest Neighbor Search
For spatial data, the nearest neighbor problem is usually solved by building a spatial index [10]. These indices can broadly be divided into two categories: those that partition the space and those that partition the data.
@@
Classic structures for space partitioning include KD Tree and quadtrees. A KD Tree [7] is a spatial partitioning structure that recursively divides the dataset along one of the k coordinate system axes, typically alternating between them; for instance, with k = 2, splits occur alternately on the x- and y-axes. This splitting process continues until a maximum depth is reached or until fewer than a specified number of points remain in a cell. KD Trees can be extended to include objects beyond points, such as triangles or line segments. Determining the optimal splitting plane can be challenging in these cases; However, for points (or disks) in two dimensions, we can simply sort the points along each dimension and split at the median. The construction of a KD Tree can be achieved in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time. A octree is another spatial partitioning structure, but it specifically divides a 3D space into eight cubes or regions. This is typically done recursively, with each subdivision occurring when the data within a region meets certain criteria, such as exceeding a maximum number of points or objects. Each internal node in the octree corresponds to a cube region of space, which is further subdivided into eight equal-sized cubes, represented by its children. The construction of an octree can be completed in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time, too.
@@
The R Tree [8] is a classic structure for partitioning data. It is based on B-trees and shares a similar structure: both the number of objects in a leaf node and the number of children in an internal node are bounded by specified minimum and maximum limits. Additionally, all leaf nodes are at the same level. The R-tree can be built in $O(n\log n)$ time and allows for the identification of the nearest neighbors in $O(\log n)$ time.
@@
Grids and linear search methods are also classic approaches for nearest-neighbor searches. The grid-based approach improves efficiency by overlaying a regular grid on the entire simulation area and tracking the agents within each cell. During a query, the search includes the cell containing the query points and the adjacent cells. The construction time for this method is $O(n+m)$, where $m$ represents the number of cells in the grid. The worst-case query time is $O(n)$, but practical performance greatly depends on the distribution of agents and the size of the cells. On the other hand, linear search is suitable when the number of query points is very small. This method finds the nearest points by calculating the distance from each target point to the query point.
II-B Voronoi Diagram & Delaunay
@@
Suppose that we have a set of generators $\mathcal{S}=\{s_{i}\}_{i=1}^{n}$ in a given domain $\Omega$ that is equipped with a metric function $\mathcal{D}$, the Voronoi diagram involves partitioning $\Omega$ into regions such that the generator $s_{i}$ dominates a region
$\text{Cell}(s_{i})=\{x\in\Omega~{}\big{|}~{}\mathcal{D}(s_{i},x)\leq\mathcal{D}(s_{j},x),j\neq i\}.$ (2)
@@
The definition of Voronoi diagrams may vary with domains, metrics, and generator types. The most prevalent version assumes that $\Omega$ represents Euclidean spaces, and the generators are exclusively points. As a fundamental tool, this has found widespread applications in computer graphics [3, 1, 11].
@@
The Delaunay is the dual structure of the Voronoi diagram, representing the adjacency relationships between the Voronoi cells. Currently, Voronoi diagrams are typically computed by first calculating the Delaunay triangulation and then deriving the dual structure. The Watson algorithm [12] is a classic method for computing the Delaunay triangulation. It is based on the empty circle property and works by incrementally inserting points and flipping triangles that do not satisfy this property. Common tools for computing Delaunay include TetGen [13] and CGAL [14], which can achieve an average computational complexity of $O(n\log n)$, with a worst-case complexity of $O(n^{2})$.
@@
Voronoi diagrams can also be applied to nearest neighbor searches. By definition, if the closest point to a query point $q$ in $\mathbf{P}$ is $p_{k}$, then $q$ lies within the Voronoi cell of $p_{k}$. This property makes Voronoi diagrams highly relevant for solving nearest neighbor problems. However, the challenge lies in efficiently identifying the specific Voronoi cell containing $q$, as there currently isn’t an efficient algorithm for this task. This limitation restricts the practical use of Voronoi diagrams in nearest neighbor searches.
II-C Farthest Point Sampling
@@
However, a notable drawback of FPS is its computational cost. The complexity of the method is $O(kn)$, where $n$ is the total number of points in the dataset, and $k$ denotes the number of points to be sampled. This complexity can become prohibitive for large datasets or when a high number of samples is needed, making FPS less efficient in such cases.
@@ III Methodology
In this section, we introduce the key concept underlying the algorithms discussed in this paper, beginning with an exploration of linear search.
We assume that the given points are numbered from 1 to $n$. Given a query point $q$, we use $\Phi_{m}(q)$ to denote the closest point in $\{p_{i}\}_{i=1}^{m}$, where $m\leq n$. This leads to the following state transfer equation:
$\Phi_{m}(q)=\begin{cases}p_{1}&\text{if }m=1,\\ \Phi_{m-1}(q)&\text{if }m\neq 1\ \&\ \|q-\Phi_{m-1}(q)\|\leq\|q-p_{m}\|,\\ p_{m}&\text{if }m\neq 1\ \&\ \|q-\Phi_{m-1}(q)\|>\|q-p_{m}\|,\end{cases}$ (4)
@@ IV-A Query Table Construction
Query Table $\mathcal{L}$ is constructed synchronously as the Voronoi diagram is incrementally built. Initially, we have only one site $p_{1}$, leaving its Query List empty. Suppose that the addition of $p_{m+1}$ changes $p_{i}$’s cell in $\mathcal{V}_{m}$. We append $p_{m+1}$ to $p_{i}$’s Query List; see Figure 2. The Query Table $\mathcal{L}$ is completed once all the sites are inserted. It is important to note that the order of elements stored in each $\mathcal{L}_{i}$ are consistent with the insertion order.
@@
Since the Delaunay is the dual of the Voronoi diagram and reflects the adjacency relationships between Voronoi cells, to check whether $p_{m}\vdash\text{Cell}(p_{i};\mathcal{V}_{m-1})$, this can be further converted into checking whether $p_{m}$ and $p_{i}$ are adjacent in the Delaunay constructed from $\{p_{i}\}_{i=1}^{m}$, which is dual to $\mathcal{V}_{m}$. Algorithm 1 provides the pseudocode for the corresponding construction process.
@@ IV-C Farthest Point Sampling
As an example, consider a state from the incremental construction of the Voronoi diagram, as shown in Figure 3(a). The black points represent the points that have already been inserted, while the gray points denote the remaining points waiting to be inserted. When a new point is inserted based on the farthest point sampling strategy, as illustrated in Figure 3(b), the red point represents the newly sampled point from the gray points, and the green area corresponds to its Voronoi cell. A key observation is that only the distances of the remaining points located within the cells adjacent to the green cell (specifically, those within these cells prior to the insertion of the red point) need to be updated. Since the white Voronoi cell does not change during the insertion of new points, this implies that the nearest neighbors of the uninserted points within it remain unchanged. This significantly reduces the computational cost of sampling. The remaining task is to dynamically track the Voronoi cells that contain the remaining points as the Voronoi diagram is incrementally constructed. This process is straightforward.
@@
IV-D Complexity Analysis
@@
$O(\sum_{k=1}^{n-1}\left[\log(n-k)+\left(c*\frac{n-k}{k}\right)\right])$ (9)
$=O(\sum_{k=1}^{n-1}\log(k)+c*n*\sum_{k=1}^{n-1}\frac{1}{k}-(n-1)*c)$ (10)
@@ V Results
In all the experiments below, Ours1 refers to our method using a random order for incremental Delaunay construction, while Ours2 indicates that the farthest point sorting strategy was applied. All the point cloud data used in the experiment has been included in the supplementary data.
@@ V-B Preprocessing Cost
When a random sorting method is adopted, the preprocessing time of the current algorithm is equivalent to that of incremental Delaunay construction. However, at this stage, our preprocessing speed is slower than that of existing Delaunay construction libraries. This is primarily due to two factors. First, our method currently supports only serial execution. Second, certain algorithms, such as TetGen, reorder points before incremental Delaunay construction to enhance overall speed. However, this insertion order is not favorable for our queries. It is worth noting, however, that our method has the potential to accelerate preprocessing when integrated with Delaunay construction. In the current implementation, we present the preprocessing times of different algorithms across several classic models in Table II. Additionally, Figure 5 shows the preprocessing time with respect to the resolution of the Dragon data, comparing our method to other methods. It can be observed that at this stage, the preprocessing time of our method exceeds that of existing nearest neighbor search algorithms.
@@
Figure 5: The preprocessing cost of KD Tree, R Tree and Ours on the Dragon data with varying resolutions. Top: preprocessing time. Bottom: the comparison about the preprocessing cost. Dashed line indicates where the ratio equals 1.
@@ V-D Tests on ABC & Thingi10K
Thingi10K [23] and ABC [24] are two large-scale 3D datasets that contain diverse models. To conduct a comprehensive comparison among KD Tree, R Tree, and Ours, we ran them on the dataset to compare both the preprocessing cost and the query cost. The query time are shown in Figures 7 and Figure 8. And the preprocessing time can be found in Figure 13 and Figure 14 in Appendix.

### v4

@@ Abstract
Given a collection of points in $\mathbb{R}^{3}$, KD-Tree and R-Tree are well-known nearest neighbor search (NNS) algorithms that rely on space partitioning and spatial indexing techniques. However, when the query point is far from the data points or the data points inherently represent a 2-manifold surface, their query performance may degrade.
To address this, we propose a novel dynamic programming technique that precomputes a Directed Acyclic Graph (DAG) to encode the proximity structure between data points. More specifically, the DAG captures how the proximity structure evolves during the incremental construction of the Voronoi diagram of the data points. Experimental results demonstrate that our method achieves a 1x-10x speedup.
Additionally, our algorithm offers several valuable features. For instance, it naturally supports an $O(k\log n)$ algorithm for farthest point sampling, where $k$ is the desired number of sample points. Moreover, density peak clustering, which involves finding the nearest point among the top K points, is typically considered to have a time complexity of $O(n^{2})$. With our algorithm, this can be reduced to $O(n\log n)$. We believe this work will inspire further research on the NNS problem.
@@ I Introduction
As shown in the inset figure, we consider two point clouds with distinct distributions: one uniformly distributed within a unit cube and the other across the surface of the Earth. Additionally, we generate two sets of query points—one within the bounding box of the input data and the other within 8 times the bounding box. It is evident that the query performance of both KD Trees and R Trees deteriorates significantly under these conditions.
@@
  • If the point cloud possesses an order, our method can facilitate the identification of the nearest point to $q$ among $\{p_{i}\}_{i=1}^{k}$, for any $k$, where $k\leq n$. This approach can be employed to reduce the time complexity of Density Peak Clustering [9], thus enhancing its applicability in large-scale scenarios.
@@
II Related Work
@@ II-A Exact Nearest Neighbor Search
For spatial data, the nearest neighbor problem is typically addressed by building a spatial indexing structure [10]. These techniques can be broadly categorized into two types: those that partition space and those that partition data.
@@
Classic structures for space partitioning include KD Trees and quadtrees. A KD Tree [7] is a spatial partitioning structure that recursively divides the dataset along one of the $k$ coordinate system axes, typically alternating between them; for instance, with $k=2$, splits occur alternately on the x- and y-axes. This splitting process continues until a maximum depth is reached or until fewer than a specified number of points remain in a cell. KD Trees can be extended to include objects beyond points, such as triangles or line segments. Determining the optimal splitting plane can be challenging in these cases; however, for points (or disks) in two dimensions, one can simply sort the points along each dimension and split at the median. The construction of a KD Tree can be achieved in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time.
An octree is another spatial partitioning structure that specifically divides 3D space into eight cubes. This is typically done recursively, with each subdivision occurring when the data within a region meets certain criteria, such as exceeding a maximum number of points or objects. Each internal node in the octree corresponds to a cube region of space, further subdivided into eight equal-sized cubes represented by its children. The construction of an octree can also be completed in $O(n\log n)$ time, and it can find the nearest neighbors in $O(\log n)$ time as well.
@@
The R Tree [8] is a classic structure for partitioning data. It is based on B-trees and shares a similar structure: both the number of objects in a leaf node and the number of children in an internal node are bounded by specified minimum and maximum limits, and all leaf nodes are at the same level. The R Tree can be built in $O(n\log n)$ time and allows for the identification of nearest neighbors in $O(\log n)$ time.
@@
Grids and linear search methods are also traditional approaches for nearest-neighbor searches. The grid-based approach improves efficiency by overlaying a regular grid on the entire simulation area and tracking agents within each cell. During a query, the search includes the cell containing the query points and the adjacent cells. The construction time for this method is $O(n+m)$, where $m$ represents the number of cells in the grid. While the worst-case query time is $O(n)$, practical performance greatly depends on the distribution of agents and the size of the cells. Conversely, linear search is suitable when the number of query points is very small. This method finds the nearest points by calculating the distance from each target point to the query point.
II-B Approximate Nearest Neighbor Search
Graph-based algorithms are highly effective for approximate nearest neighbor (ANN) search due to their balance of accuracy and computational efficiency [11, 12, 13, 14]. Typically, the process begins with constructing connections within a given point cloud, followed by iterative traversal of the graph starting from an entry point, which may be selected randomly or determined by a separate algorithm. At each step, the algorithm selects the closest neighboring node as the next base node, iteratively refining the current solution. The search terminates once a predefined stopping condition, such as a set number of distance evaluations, is met.
One well-known graph structure used for search is the Navigable Small World (NSW, also known as Metricized Small World, MSW) [15, 14, 12]. NSW constructs the graph by inserting elements in random order. Each element is then bidirectionally connected to its $M$ nearest neighbors among the previously inserted nodes. To improve search efficiency, many algorithms adopt a hierarchical strategy by creating multiple layers of the graph, which helps narrow the search space and accelerate the process.
Our method adopts a similar graph-based strategy, emphasizing efficient connectivity and traversal. Unlike traditional graph-based approximate nearest neighbor search, our method precisely identifies the nearest points with strict guarantees of correctness. Furthermore, although we do not explicitly use hierarchical structures, our integrated construction and traversal strategies achieve comparable benefits, such as efficient narrowing of the search space and improved accuracy.
II-C Voronoi Diagram & Delaunay
@@
Suppose we have a set of generators $\mathcal{S}=\{s_{i}\}_{i=1}^{n}$ in a given domain $\Omega$ equipped with a metric function $\mathcal{D}$. The Voronoi diagram partitions $\Omega$ into regions such that the generator $s_{i}$ dominates a region defined as
$\text{Cell}(s_{i})=\{x\in\Omega~{}|~{}\mathcal{D}(s_{i},x)\leq\mathcal{D}(s_{j},x),\,\forall j\neq i\}.$
@@
The definition of Voronoi diagrams may vary based on domains, metrics, and types of generators. The most prevalent version assumes that $\Omega$ represents Euclidean spaces and that the generators are exclusively points. As a fundamental tool, Voronoi diagrams have found widespread applications in computer graphics [3, 1, 16].
@@
The Delaunay triangulation is the dual structure of the Voronoi diagram, representing the adjacency relationships between the Voronoi cells. Currently, Voronoi diagrams are typically computed by first calculating the Delaunay triangulation and then deriving the dual structure. The Watson algorithm [17] is a classic method for computing the Delaunay triangulation. It is based on the empty circle property and works by incrementally inserting points and flipping triangles that do not satisfy this property. Common tools for computing Delaunay triangulations include TetGen [18] and CGAL [19], which can achieve an average computational complexity of $O(n\log n)$ with a worst-case complexity of $O(n^{2})$.
@@
Voronoi diagrams can also be applied to nearest neighbor searches. By definition, if the closest point to a query point $q$ in $\mathbf{P}$ is $p_{k}$, then $q$ lies within the Voronoi cell of $p_{k}$. This property makes Voronoi diagrams highly relevant for solving nearest neighbor problems. However, there is currently no efficient algorithm for identifying the specific Voronoi cell containing $q$. This restricts the practical use of Voronoi diagrams in nearest neighbor searches.
II-D Farthest Point Sampling
@@
The time complexity of a typical FPS method [20, 21] is $O(kn)$, where $n$ is the total number of points in the dataset, and $k$ denotes the number of points to be sampled. The latest FPS method QuickFPS [22] organizes large-scale point clouds into multiple bins. By employing two mechanisms—merge computation and implicit computation on these bins—it significantly reduces external memory access and computation costs, thereby improving the speed of farthest point sampling.
@@ III Methodology
In this section, we introduce the key concept underlying the algorithms, beginning with a solution of linear search.
We assume that the given points are numbered from 1 to $n$. Given a query point $q$, we use $\Phi_{m}(q)$ to denote the closest point in the subset $\{p_{i}\}_{i=1}^{m}$, where $m\leq n$. This leads to the following state transfer equation:
$\Phi_{m}(q)=\begin{cases}p_{1}&\text{if }m=1,\\ \Phi_{m-1}(q)&\text{if }m\neq 1\land\|q-\Phi_{m-1}(q)\|\leq\|q-p_{m}\|,\\ p_{m}&\text{if }m\neq 1\land\|q-\Phi_{m-1}(q)\|>\|q-p_{m}\|,\end{cases}$ (3)
@@ IV-A Query Table Construction
$\mathcal{L}$ is constructed simutaneously as the Voronoi diagram is incrementally built. Initially, we have only one site $p_{1}$, leaving its Query List empty. Suppose that the addition of $p_{m+1}$ changes $p_{i}$’s cell in $\mathcal{V}_{m}$. We append $p_{m+1}$ to $p_{i}$’s Query List; see Figure 2. The Query Table $\mathcal{L}$ is completed once all the sites are inserted. It is important to note that the order of elements stored in each $\mathcal{L}_{i}$ is consistent with the insertion order.
@@
Since the Delaunay is the dual of the Voronoi diagram and encodes the adjacency relationships between Voronoi cells, to check whether $p_{m}\vdash\text{Cell}(p_{i};\mathcal{V}_{m-1})$, this can be further converted into checking whether $p_{m}$ and $p_{i}$ are adjacent in the Delaunay constructed from $\{p_{i}\}_{i=1}^{m}$, which is dual to $\mathcal{V}_{m}$. Algorithm 1 provides the pseudocode for the corresponding construction process.
@@ IV-C Farthest Point Sampling
As an example, consider a state from the incremental construction of the Voronoi diagram, as shown in Figure 3(a). The black points represent the points that have already been inserted, while the gray points denote the remaining points waiting to be inserted. When a new point is inserted based on the farthest point sampling strategy, as illustrated in Figure 3(b), the red point represents the newly sampled point from the gray points, and the green area corresponds to its Voronoi cell. A key observation is that only the distances of the remaining points located within the cells adjacent to the green cell (specifically, those within these cells prior to the insertion of the red point) need to be updated. Since the white Voronoi cell does not change during the insertion of new points, this implies that the nearest neighbors of the uninserted points within it remain unchanged. This significantly reduces the computational cost of sampling. The remaining task involves two parts: dynamically tracking the remaining points contained within the cells of the already inserted points during the incremental construction of the Voronoi diagram, and dynamically obtaining the next sampling point. The first part can be accomplished efficiently due to the local properties of Voronoi cells, while for the second part, we utilize a segment tree structure in our implementation.
@@
IV-D Incremental Delaunay Construction
Currently, numerous incremental methods for Delaunay construction, such as parallel acceleration [24, 25] and pre-sorting points [26], are commonly used to speed up the process. However, due to the specific requirements of farthest point sorting and the need to generate a query table during the construction phase in our approach, traditional acceleration techniques for Delaunay construction are not directly applicable.
During the incremental Delaunay construction process, one of the most time-consuming operations is locating the tetrahedron that contains a newly inserted point. According to Section IV-A and IV-B of our algorithm, when inserting point $p_{k}$, we have already constructed a query table for finding the nearest point within the set $\{p_{i}\}_{i=1}^{k-1}$. This query table stores information that allows for rapid identification of nearest points among the already inserted points. Leveraging this structure, we can quickly determine the nearest point to $p_{k}$. This nearest point then serves as the starting point for efficiently locating the tetrahedron containing $p_{k}$, significantly accelerating the Delaunay construction process.
IV-E Complexity Analysis
@@
$O(\sum_{k=1}^{n-1}\left[\log(n-k)+\left(c\times\frac{n-k}{k}\right)\right])$
$=O(\sum_{k=1}^{n-1}\log(k)+cn\times\sum_{k=1}^{n-1}\frac{1}{k}-(n-1)\times c)$
@@ V Results
In all the experiments below, Ours1 refers to our method using a random order for incremental Delaunay construction, while Ours2 indicates that the farthest point sorting strategy was applied.
@@ V-B Preprocessing Cost
The preprocessing time of our algorithm is primarily concentrated in the Delaunay construction phase, as it involves significant computational overhead in building and managing the complex graph structure. In this section, we present the preprocessing times of different algorithms across several classic models in Table II. Additionally, Figure 5 illustrates the preprocessing time relative to the resolution of the Dragon point cloud, providing a comparative analysis between our method and others. Furthermore, Figures 9 and Figure 10 show the preprocessing time of different methods on subsets of the ABC dataset [33] and the Thingi10k dataset [34]. It can be observed that at this stage, our method currently exhibits longer preprocessing times compared to existing nearest neighbor search algorithms, particularly at higher resolutions, indicating areas that require further optimization.
@@
Figure 5: The preprocessing cost of KD Tree, R Tree and Ours on the Dragon data with varying resolutions. Top: preprocessing time. Bottom: the comparison about the preprocessing cost.
[…]
@@ V-D Tests on ABC & Thingi10K
Thingi10K [34] and ABC [33] are two large-scale 3D datasets that contain diverse models. To conduct a comprehensive comparison among KD Tree, R Tree, and Ours, we ran them on the dataset to compare both the preprocessing cost and the query cost. The query time are shown in Figures 9 and Figure 10. And the preprocessing time can be found in Figure 6 and Figure 7.
@@ V-E Other Types of Data
V-G Farthest point sampling
QuickFPS [22] is currently the fastest farthest point sampling strategy, achieving performance speedups of 230.5× over traditional CPU implementations, 43.4× over traditional GPU implementations, and 12.2× over previous state-of-the-art point cloud accelerators. It divides the point cloud into multiple buckets based on spatial location, and utilizes advanced techniques such as merged computation and implicit computation to accelerate the farthest point sampling process. Although the method proposed in this paper is not specifically designed for farthest point sampling, it also demonstrates impressive speed in this task. Using several traditional models as examples, with the farthest point sampling rate set at 25%, the sampling times of QuickFPS (CPU Version) and the method proposed in this paper are shown in Table IV. Moreover, compared to the complexity of QuickFPS, our proposed algorithm are both simpler and more intuitive.
TABLE IV: Comparison of Sampling Times (s) between QuickFPS and Ours at a 25% Sampling Rate.
Camel Bunny Dragon Kitten Armadillo Lucy Sponza
Vertices 28934 72911 100313 291023 726367 1018219 1313504
QuickFPS 0.016 0.094 0.156 1.109 2.391 2.875 3.969
Ours 0.079 0.25 0.328 1.484 5.421 8.453 11.25